4 June 1999, evening shift summary

Nickolas Solomey (solomey@hep.uchicago.edu)
Fri, 4 Jun 1999 23:57:21 -0500

Shift Summary for evening shift, 06/04/99

On Shift: Sandrine and Nick

Other present: Jason 1, Jason 2, Ed, Mike A., David N., Sydney, Rick C.,
Rick F., Angela, Patty, Migual, Hogan, Ed, Roland, Herman, Brian L.,
Sasha, Tsuyoshi, Jiambo, and Elizabeth

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity (1.2E12): 0
Hours of E799 data: 0
Hours of special runs: none except laser scans, pedistal, and cosmic
without writing to tape.

Accounting of lost time: KTeV on access for the next day.

Runs ended this shift:

13358: Laser Scan
13359: Cosmic run - garbage run
13361: Cosmic run - in order to monitor channel 2344 (timing problem)
13362: KTeV_PED -
13366: long Cosmic run to keep everything working.

Accesses during shift: Open access for all shift.

Description of shift:

15:30: Came on shift and found that there was work going on in
the area of DC4, DC2. Vivian was working with software repairs of
the DAQ system.

16:00: Jason H. replaced the ET comparator of channel 1534 since there
was a hot bit.

16:15: Rick C. came back from target hall, said he changed the PMT ACC1-2
for the accidential counter and put back 3 fold coincidence. We still
have 1/8" x 1/8" hole (last run was 1/2" x 1/2")

19:35: Ed and Sandrine swapped the DPMT of channel 2334. The new gain for
this channel has shifted by 3.5 %, so we don't change the gain until the
next calibration. Start a pedestal run.

23:50 A quiet shift, wish we had more excitment, next time hope
for beam and lots of data to tape. We are almost there!

Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey